# Query-Guided Self-Supervised Summarization of Nursing Notes

### v1

@@ Abstract
Nursing notes, an important component of Electronic Health Records (EHRs), keep track of the progression of a patient’s health status during a care episode. Distilling the key information in nursing notes through text summarization techniques can improve clinicians’ efficiency in understanding patients’ conditions when reviewing nursing notes. However, existing abstractive summarization methods in the clinical setting have often overlooked nursing notes and require the creation of reference summaries for supervision signals, which is time-consuming. In this work, we introduce QGSumm, a query-guided self-supervised domain adaptation framework for nursing note summarization. Using patient-related clinical queries as guidance, our approach generates high-quality, patient-centered summaries without relying on reference summaries for training. Through automatic and manual evaluation by an expert clinician, we demonstrate the strengths of our approach compared to the state-of-the-art Large Language Models (LLMs) in both zero-shot and few-shot settings. Ultimately, our approach provides a new perspective on conditional text summarization, tailored to the specific interests of clinical personnel.
@@ 1 Introduction
Electronic Health Records (EHRs) document the events that patients go through during their hospitalization. These records consist of both free-text clinical notes and structured data. Among them, nursing notes are important for keeping track of the progression of patients’ illnesses, changes in health status, as well as the medications and procedures administered (Törnvall and Wilhelmsson, 2008). Nursing notes provide clinicians with comprehensive insights into the patients’ conditions, assisting in formulating next-step treatment and care plans, as well as in writing the final discharge summaries. However, a patient’s care episode may result in a large number of nursing notes, especially for patients suffering from complex health problems, which causes the problem of information overload (Hall and Walton, 2004). Additionally, the information in nursing notes is usually intricate and highly condensed, making it time-consuming for clinicians to understand (Clarke et al., 2013).
In Natural Language Processing (NLP), text summarization techniques can be used to distill the content of nursing notes (Wang et al., 2021) to help clinicians quickly grasp their contents. Automatic clinical note summarization has been extensively studied, with existing approaches categorized into extractive (Pivovarov and Elhadad, 2015; Moen et al., 2016; Tang et al., 2019) and abstractive methods (Zhang et al., 2020b; Liu et al., 2022a; Searle et al., 2023). However, these methods have certain limitations: (1) Extractive methods only retain sentences, important words, or keyphrases from the original note, limiting coherence and fluency of the summary. This presents a challenge for understanding the summarized content. On the other hand, (2) abstractive methods can generate smoother summaries, but most of the existing works on abstractive clinical note summarization require explicit supervision, i.e., a reference summary as the ground truth. Writing the references is time-consuming (O’Donnell et al., 2009), causing a shortage of training data. Moreover, (3) most abstractive clinical note summarization methods focus on a specific type of notes, such as discharge summaries, radiology reports, or the dialogues between doctors and patients, and there is a lack of research on abstractive nursing note summarization.
@@
To address these limitations, we propose a novel approach for abstractive nursing note summarization, which does not require reference summaries for training. The nature of the nursing documentation poses additional challenges. For example, as shown in Figure 1, information in nursing notes may lack clarity, be poorly organized, and contain medical jargon that often includes non-standard abbreviations. In the absence of supervised learning signals, guiding the summarization model to understand the semantic information in notes and generate a good summary becomes challenging. Some previous text summarization works (Chu and Liu, 2019; Elsahar et al., 2021) adapt strategies in self-supervised learning (Liu et al., 2021) to such a scenario. In their methods, the training objective is to decrease the semantic distance between a summary and the original text based on the assumption that a good summary is capable of reconstructing the source text. However, simply making the semantic representation of the summary close to that of the original document does not allow controlling the generated summaries, which may result in a lack of relevant information. In clinical domains, we have specific requirements for the content of the summary, where the focus should be on the patient’s condition. Thus, methods that rely solely on the semantic similarity may not be fully satisfactory.
A good summary of a clinical note is centered on the patient’s condition. Consequently, when queried about the state of the patient, answers obtained from the summary will be similar to those obtained from the original note.
For instance, a query can concern the probability of a patient’s condition improving in the near future. We can train a model to answer this query using the current nursing note or its summary as input, and if the summary is accurate these two answers should be similar. Accordingly, we design a learning objective that aims at minimizing the discrepancy between the responses from the summary or from the source note to given queries. To further encourage the model to prioritize the patient’s current medical condition, we integrate into the summarization workflow both the patient’s metadata as well as information in the previous notes of the patient recorded on the same admission.
To the best of our knowledge, our proposed framework is the first on abstractive summarization of nursing notes, and there is no previous work on employing the self-supervised learning strategy for clinical note summarization. Our primary contributions are:
  • The study focuses on nursing notes that play a critical role in clinical settings, filling a gap in previous research by introducing a method for abstractive nursing note summarization. Our method’s ability to operate effectively without requiring reference summaries highlights its practical applicability.
  • We propose a novel self-supervised domain adaptation framework. By leveraging patient-related queries to guide the model, we achieve the goal of generating nursing note summaries that prioritize specific content, i.e., patients’ conditions and health status, without the need for ground truth.
  • We conduct a comprehensive automatic evaluation and a manual evaluation by an expert clinician. We compare our approach with state-of-the-art Large Language Models (LLMs) in few- and zero-shot settings. This demonstrates the method’s ability to generate high-quality summaries of nursing notes, and additionally provides an independent evaluation of the common LLMs in this task.
Generalizable Insights about Machine Learning in the Context of Healthcare
In existing healthcare-related NLP research, there is a noticeable gap in addressing nursing notes specifically. Summarizing key patient information within nursing notes could enhance the efficiency of medical personnel’s workflow. We provide a new perspective on obtaining summaries of nursing notes through self-supervised domain adaptation without the need for manually written reference summaries. In unconditional text summarization, the generated summary lacks explicit constraints. On the other hand, most conditional text summarization methods typically require data annotation (Vig et al., 2022) or the extraction of information related to content conditions (Pagnoni et al., 2023) from the source text, making them less suitable for our task. We introduce easily applicable patient-related queries as a way to facilitate conditional text summarization of nursing notes, which ensure that the generated summaries contain information required to respond to the query effectively, and are closely linked to the patient’s key information. Such constraints and guidance make our method highly suitable for the clinical and healthcare field since the resulting summaries are centered on the information nurses and clinicians are most concerned about.
@@ 2 Related Work
Key Information Extraction and Summarization from Clinical Notes
Research in this domain focuses on two approaches, extractive and abstractive summarization. The extractive method can preserve faithfulness but results in the inability to paraphrase and difficulties in maintaining coherence. Earlier work primarily used semantic similarity-based techniques (Pivovarov and Elhadad, 2015; Moen et al., 2016). The emergence of Transformer models has shifted the focus to attention-based methods to determine key information in clinical text with an emphasis on explainability (Tang et al., 2019; Reunamo et al., 2022; Kanwal and Rizzo, 2022).
Recently, there has been a notable increase in research on abstractive clinical text summarization. From an application perspective, these methods mainly target discharge summary generation (Shing et al., 2021; Adams et al., 2022; Searle et al., 2023), radiology report summarization (Zhang et al., 2020b; Van Veen et al., 2023), summarization of doctor-patient conversations (Zhang et al., 2021; Krishna et al., 2021; Abacha et al., 2023) and problem list summarization (Gao et al., 2022, 2023). However, unlike our approach, most of these works depend on data annotation or reference summaries for training and domain adaptation.
LLMs demonstrate a remarkable capability in clinical text understanding, leading to interest in investigating their performance in clinical text summarization. Van Veen et al. (2024) extensively analyze the clinical text summarization performance of various LLMs with in-context learning (Lampinen et al., 2022) and QLoRA (Dettmers et al., 2024) adaptation. They compare the performance of LLMs with medical experts, providing insights into the strengths and limitations of LLMs.
Unsupervised and Self-Supervised Abstractive Text Summarization
The scarcity of annotated text for abstractive text summarization tasks has spurred interest in unsupervised and self-supervised text summarization. Previous works have relied on source document reconstruction, operating under the assumption that a good summary should be able to reconstruct the source document or capture its essential content (Chu and Liu, 2019). However, reconstructing an entire text using a summary without any guiding signal or prompt is challenging. In contrast, Yang et al. (2020) leverage the lead bias in news articles, by pretraining a model to predict the leading sentences as the target summary. However, this approach requires specific information distribution and text layout, which is not generally applicable. Some works have proposed two-step approaches to first extract important information or entities in the source text and then perform abstractive summarization with the guidance of this extracted information (Zhong et al., 2022; Ke et al., 2022; Liu et al., 2022b). However, the quality of the generated summary relies on the effectiveness of the extraction process, and developing a reliable extractor may entail significant costs. Zhuang et al. (2022) propose a contrastive learning strategy, using source documents as positive and edited source documents as negative examples. Their training objective aims at maximizing the semantic similarity between generated summaries and positive examples while minimizing those between generated summaries and negative examples. Hosking et al. (2023) propose an attributable opinion summarization system, which encodes sentences as paths through a hierarchical discrete latent space. Given a specific entity, the system can identify its common subpaths that are decoded as the output summary.
@@ 3 Methods
Figure 2: The overall architecture. While fine-tuning the encoder (ENC) and decoder (DEC), $\operatorname{DEC^{rec}}$ and the query responder network $\operatorname{R}$ are frozen. Raw patient information is processed by ENC into the embeddings $\mathbf{H}^{PA}$, but this is omitted here for clarity (see text and Fig. 3 for details).
Next, we introduce QGSumm, a novel framework to automatically summarize and refine clinical notes, with a focus on capturing important patient-centered information in a self-supervised fashion (Figure 2). In line with the hypothesis, we propose a self-supervised domain adaptation strategy applied on the base model presented in Section 3.1. This strategy positive-contrastively learns from the original nursing notes, providing the summaries with an ability comparable to the original notes to resolve patient-related queries (Section 3.2). Moreover, we aim for the model to maintain focus on the patient’s meta information while also considering temporal aspects during the generation process. To achieve this, we propose two augmentation blocks, detailed in Section 3.3, to enhance the overall performance. Our model summarizes one nursing note at a time, taking into account its context. Assume a patient $PT$ has a sequence of nursing notes $N=\{N_{1},N_{2},\ldots,N_{m}\}$ sorted by time. Our objective is to obtain a summary $S_{i}$ for note $N_{i}$ from the distribution $P(S_{i}|N_{i},{PA},\{N_{1},\ldots,N_{i-1}\},U)$, which is conditioned on the patient’s metadata ${PA}$, information in the past notes $\{N_{1},\ldots,N_{i-1}\}$, and the user query $U$ guiding the generation.
@@ 3.1 Base Model
The backbone of our framework is an off-the-shelf transformer-based language model with an encoder-decoder structure, denoted by $\operatorname{M}$. Specifically, we leverage a checkpoint $\operatorname{M^{sum}}$ of $\operatorname{M}$ as the base model, which has been fine-tuned for text summarization using publicly available datasets. This allows us to efficiently utilize the extensive resources offered by the pre-trained language model without the effort of training from scratch. Hence, $\operatorname{M^{sum}}$ has been enriched with task-specific knowledge for improved performance in text summarization. However, the capability of $\operatorname{M^{sum}}$ to understand clinical text still remains limited. Therefore, additional refinement of $\operatorname{M^{sum}}$ is necessary to enhance its ability to grasp the complicated semantic information within nursing notes.
Let $N_{i}=[t_{1},t_{2},....,t_{n}]$, where $t_{i},i=1,\ldots,n$, denote tokens in $N_{i}$. As a preliminary step, we first train the encoder $\operatorname{ENC}$ of the base model $\operatorname{M^{sum}}$ by reconstructing $N_{i}$:
$\mathcal{L}_{rec}(N_{i},\operatorname{ENC},\operatorname{DEC^{rec}})=\mathcal{L}_{CrossEntropy}(N_{i},\operatorname{DEC^{rec}}(\operatorname{ENC}(N_{i})).$ (1)
Here, $\operatorname{DEC^{rec}}$ is the decoder of the original pretrained model $\operatorname{M}$, which remains frozen during the training. This process empowers the encoder with the ability to understand the semantic information and the clinical knowledge embedded in nursing notes, enabling it to encode the notes more effectively. This preparatory step as precedes the primary workflow for nursing notes summarization.
@@ 3.2 Training Objective
Since there is no ground truth summary available, the conventional method to guide the model $\operatorname{M^{sum}}$ through supervised fine-tuning is not feasible. Instead, we adopt a self-supervised strategy to force the model to generate high-quality, patient-centered summaries that can respond to patient-related queries effectively. We introduce a model $\operatorname{R}$, which serves as a query responder. This model has been trained to generate responses to specific queries concerning the patient. For example, if a query pertains to the patient’s readmission status, $\operatorname{R}$ is trained to classify patients based on readmission risk using data available in the patient database.
When giving the original nursing note $N_{i}$ or its summary $S_{i}$ generated by $\operatorname{M^{sum}}$ as an input to the responder $\operatorname{R}$, the training objective is to minimize the discrepancy between the two responses:
$\operatorname{min}\mathcal{L}_{CrossEntropy}(\operatorname{R}(N_{i}),\operatorname{R}(S_{i})).$ (2)
This formulation ensures that when responding to a certain patient-related query, using the summary will result in a response similar to that obtained using the original nursing note. To prevent $\operatorname{M^{sum}}$ from generating summaries that are too verbose or direct “copy-paste” from the original notes, we introduce a length penalty term. Therefore, the final loss function for nursing notes within one batch becomes:
$\mathcal{L}_{summ}=\frac{1}{K}\sum_{r=1}^{K}\mathcal{L}_{CrossEntropy}(\operatorname{R}(N_{r}),\operatorname{R}(S_{r}))\times(1+\lambda_{1}e^{(\alpha-0.5)}),$ (3)
@@
$\alpha=\frac{\sum_{r=1}^{K}\operatorname{Len}(S_{r})}{\sum_{r=1}^{K}\operatorname{Len}(N_{r})},$ (4)
$K$ is the batch size, and $\operatorname{Len}$ denotes the length of the document in terms of the number of tokens. The hyperparameter $\lambda_{1}\in[0,1]$ regulates the extent of the penalty. The information flow from $\operatorname{M^{sum}}$ and $\operatorname{R}$ introduces nondifferentiability into the framework, and we resolve it using the straight-through gumbel softmax trick (Bengio et al., 2013; Jang et al., 2017).
@@ 3.3 Augmentation Blocks for the Context of the Patient
A patient typically has multiple nursing notes ordered in time to document the evolution of her condition. Therefore, the key information crucial for summarizing a patient’s current status is influenced by the context provided by the prior notes. We regard this as temporal information which should be incorporated during summarization to help the model understand the progression of the patient’s condition.
For $N_{i}$, the embeddings of its previous notes are represented by the embeddings of their respective first tokens, which are special tokens indicating the start of each note. These embeddings are obtained at the last hidden state in the $\operatorname{ENC}$, denoted as $\{\mathbf{h_{1}},\mathbf{h_{2}},...,\mathbf{h_{i-1}}\}$, where $\mathbf{h_{i}}\in\mathbb{R}^{d}$ and $d$ is the dimension of the hidden state. We aggregate the representations of the past notes by weighted mean pooling such that the most recent notes receive the largest weight. In practice, we determine initial weights $\beta_{j},j=1,\ldots,i-1$ for each past note $N_{i}$ based on the position in the sequence, such that $\beta_{1}=1,\beta_{2}=2,\ldots,\beta_{i-1}=i-1$. The weighted mean pooling is performed using normalized weights:
$\beta^{{}^{\prime}}_{j}=\frac{\beta_{j}}{\beta_{1}+\beta_{2}+...+\beta_{i-1}},$ (5)
@@
where $\mathbf{h}^{TIF}\in\mathbb{R}^{d}$ represents the information fusion of the past notes. As shown in Figure 3, we prepend a special token [TI] at the beginning of the decoder input, representing temporal information with embedding $\mathbf{h}^{TIF}$. Consequently, the initial input to the decoder at the first time step consists of [[TI], [BOS]], where [BOS] is a special token indicating the start of generation. We substitute [TI] with the padding token [PAD] for nursing notes that have no past notes.
The model generates subsequent tokens in the summary in an auto-regressive manner. At each time step, the token produced is appended at the end of the decoder input for generation of subsequent tokens. Therefore, the [TI] token contributes to the generation of each token in the summary, serving as a prompt which consistently guides the model to focus on information about the patient’s past condition.
@@
We aim at obtaining summaries focusing on the patient’s condition. To aid this, we explicitly incorporate patient-level information into the model through a cross-attention mechanism, which facilitates the interaction of information on different levels of representation learning. A patient’s metadata PA typically comprise basic information recorded for the patient’s admission, including age, gender, existing diagnoses, and performed procedures. We convert this metadata into patient information in natural language (one example presented in Appendix A.1), and then encode it using $\operatorname{ENC}$ to derive patient embedding $\mathbf{H}^{PA}\in\mathbb{R}^{z\times d}$ for patient $PT$, where $z$ represents the number of tokens in patient information. The encoder also learns the embedding of the source note, $\mathbf{H}^{enc}\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of tokens in the note given as input to the encoder. On the decoder $\operatorname{DEC}$ side, let us assume the tokens input to the decoder at the current timestep are [[TI], [BOS], $y_{1}$,…,$y_{j}$]. Consequently, the hidden representation passed to the $l$th decoder layer is $\mathbf{H}_{l}^{dec}\in\mathbb{R}^{{(j+2)}\times d}$. The hidden representation $\mathbf{H}_{l}^{dec}$ is processed and updated in each decoder layer using the conventional self-attention and cross-attention with $\mathbf{H}^{enc}$. Furthermore, we augment the decoder layer with patient information by performing cross-attention also between the hidden representation $\mathbf{H}_{l}^{dec}$ and the patient embedding $\mathbf{H}^{PA}$. This facilitates the fusion of patient- and note-level information:
$\mathbf{H}_{l+1}^{dec}=\operatorname{MHCA}(\mathbf{H}^{enc},\operatorname{MHSA}(\mathbf{H}_{l}^{dec}))+\lambda_{2}\times\operatorname{MHCA}(\mathbf{H}^{PA},\operatorname{MHSA}(\mathbf{H}_{l}^{dec})),$ (7)
@@
where the $\operatorname{MHCA}$ and the $\operatorname{MHSA}$ respectively denote Multi-Head Cross-Attention and Multi-Head Self-Attention (Vaswani et al., 2017). $\lambda_{2}\in[0,1]$ is a hyperparameter to control the importance of patient meta information. $\mathbf{H}_{l+1}^{dec}$ is the input to the next decoder layer, or if the $l$th layer is the final layer, it is the input to the language modeling head.

### v2

@@ Abstract
Nursing notes, an important part of Electronic Health Records (EHRs), track a patient’s health during a care episode. Summarizing key information in nursing notes can help clinicians quickly understand patients’ conditions. However, existing summarization methods in the clinical setting, especially abstractive methods, have overlooked nursing notes and require reference summaries for training. We introduce QGSumm, a novel query-guided self-supervised domain adaptation approach for abstractive nursing note summarization. The method uses patient-related clinical queries for guidance, and hence does not need reference summaries for training. Through automatic experiments and manual evaluation by an expert clinician, we study our approach and other state-of-the-art Large Language Models (LLMs) for nursing note summarization. Our experiments show: 1) GPT-4 is competitive in maintaining information in the original nursing notes, 2) QGSumm can generate high-quality summaries with a good balance between recall of the original content and hallucination rate lower than other top methods. Ultimately, our work offers a new perspective on conditional text summarization, tailored to clinical applications.
keywords:
abstractive text summarization, nursing notes, self-supervised learning
@@ 1 Introduction
Nursing notes are important for clinicians to track a patient’s health status and administered treatments during hospitalization (Törnvall and Wilhelmsson, 2008). However, a care episode may result in a large number of nursing notes, especially for patients suffering from complex health problems (Hall and Walton, 2004). Furthermore, the condensed nature of nursing notes makes them time-consuming for clinicians to understand (Clarke et al., 2013). Text summarization methods from Natural Language Processing (NLP) help distill the content of clinical notes (Wang et al., 2021), making crucial information more accessible and less time-consuming to review. This can be particularly beneficial for clinicians, and for conducting retrospective studies. These summarization techniques are divided into extractive (Pivovarov and Elhadad, 2015; Moen et al., 2016; Tang et al., 2019) and abstractive (Zhang et al., 2020b; Liu et al., 2022a; Searle et al., 2023). Extractive methods retain the original sentences and/or keyphrases but may lack coherence or fluency. On the other hand, abstractive methods produce smoother summaries but typically require explicit supervision, i.e., a reference summary as the ground truth, which are time-consuming to produce(O’Donnell et al., 2009).
[…]
fig:example Refer to caption
@@
Self-supervised abstractive text summarization (Chu and Liu, 2019; Elsahar et al., 2021) bypasses the problem of lacking ground-truth summaries. Previous self-supervised methods (e.g., Liu et al. (2021)) reduce the semantic distance between a summary and the original text. However, unconditionally making the semantic representations similar lacks control of the generated summaries and may result in missing relevant information, which can be important especially in clinical applications. Furthermore, the content of nursing notes may be unclear, poorly organized and contain non-standard abbreviations (Figure LABEL:fig:example). Hence, in the absence of supervised learning signals, guiding the model to generate a good summary becomes challenging.
In this paper, we propose a query-guided self-supervised domain adaptation framework for nursing note summarization, named QGSumm. A good summary of a clinical note is centered on the patient’s condition.
We achieve this by designing queries that concern the patient’s condition, and we train the model to produce summaries such that answers to the queries based on the summaries are similar to the answers based on the original notes. This guidance makes our method highly suitable for the clinical field since the resulting summaries are centered on the information nurses and clinicians are most concerned about. To the best of our knowledge, our study is the first on abstractive summarization of nursing notes and on employing self-supervised learning for clinical note summarization.
Our primary contributions are: (1) The study focuses on nursing notes that play a critical role in clinical settings, filling a gap in previous research by introducing a method for abstractive nursing note summarization. Our method’s ability to work without reference summaries highlights its practical applicability, while also being relatively light-weight and suitable for deployment in a secure environment with limited computing resources. (2) We propose a novel self-supervised domain adaptation framework. By leveraging patient-related queries, we guide the model to generate nursing note summaries that prioritize specific content, i.e., the patients’ conditions and health status, without the need for manually written reference summaries as ground truth. (3) We conduct a comprehensive automated empirical study and a manual evaluation by an expert clinician, including state-of-the-art Large Language Models (LLMs), which have not been previously investigated for nursing note summarization. The experiments demonstrate especially our method’s and GPT-4’s ability to perform well in this task.
[…]
fig:framework Refer to caption
Figure 2: Stage 1: Fine-tune the encoder ($\operatorname{ENC_{0}}$) by note reconstruction, where $\operatorname{DEC_{0}}$ is frozen; Stage 2: Fine-tune the encoder (ENC) and decoder (DEC) in the self-supervised manner, where the query responder $\operatorname{R}$ is frozen. Raw patient information is processed by ENC into the embeddings $\mathbf{H}^{PA}$, as shown in Fig. LABEL:fig:augmentation
@@ 2 Related Work
Clinical Notes Summarization.
Extractive summarization methods can preserve faithfulness but may have difficulties in maintaining coherence, and include earlier semantic similarity-based techniques (Pivovarov and Elhadad, 2015; Moen et al., 2016), as well as more recent attention-based methods (Tang et al., 2019; Reunamo et al., 2022; Kanwal and Rizzo, 2022). On the other hand, abstractive clinical notes summarization has been applied on discharge summaries (Shing et al., 2021; Adams et al., 2022; Searle et al., 2023), radiology reports (Zhang et al., 2020b; Van Veen et al., 2023; Nishio et al., 2024), doctor-patient conversations (Zhang et al., 2021; Krishna et al., 2021; Abacha et al., 2023) and problem lists (Gao et al., 2022, 2023). However, unlike ours, these works depend on data annotation or reference summaries. LLMs demonstrate a remarkable capability in clinical text understanding. Van Veen et al. (2024) extensively analyze the clinical text summarization performance of various LLMs with in-context learning (Lampinen et al., 2022) and QLoRA (Dettmers et al., 2024) adaptation.
Unsupervised and Self-Supervised Abstractive Text Summarization.
The scarcity of annotated text has spurred interest in unsupervised and self-supervised text summarization. Previous works like source document reconstruction assume that a good summary can reconstruct the source document (Chu and Liu, 2019), but such summaries may lack specific information due to limited control. Two-step approaches extract important information first and then perform abstractive summarization based on this extraction (Zhong et al., 2022; Ke et al., 2022; Liu et al., 2022b). However, the quality of such summaries relies on the extraction effectiveness, and developing a reliable extractor can be costly. A contrastive learning strategy, proposed by Zhuang et al. (2022), aims to maximize similarity between generated summaries and source documents while minimizing it between summaries and edited documents. Hosking et al. (2023) suggest an attributable opinion summarization system that encodes sentences as paths through a hierarchical discrete latent space, identifying common subpaths for an entity to generate the summary. Jin and Chen (2024) propose using a review from a review set as the hypothetic summary to carry out self-supervised summarization of product reviews.
@@ 3 Methods
Next, we introduce QGSumm, a novel framework for summarizing clinical notes, with a focus on capturing important patient-centered information in a self-supervised fashion. We propose a self-supervised domain adaptation strategy applied on the base model in Section 3.1. This strategy positive-contrastively learns from the original nursing notes, providing the summaries with an ability comparable to the original notes to resolve patient-related queries (Section 3.2). Using two augmentation blocks (Section 3.3), the model leverages patient metadata and temporal aspects. Assume a patient $PT$ has a sequence of nursing notes $N=\{N_{1},N_{2},\ldots,N_{m}\}$ sorted by time. Our objective is to obtain a summary $S_{i}$ for note $N_{i}$ from the distribution $P(S_{i}|N_{i},{PA},\{N_{1},\ldots,N_{i-1}\},U)$, which is conditioned on the patient’s metadata ${PA}$, information in the previous notes $\{N_{1},\ldots,N_{i-1}\}$, and the query $U$ which helps guide the generation.
@@ 3.1 Base Model
[…]
fig:augmentation Refer to caption
The backbone of our framework is an off-the-shelf transformer-based language model with an encoder-decoder structure. We leverage its checkpoint $\operatorname{M_{0}}$, fine-tuned for text summarization, as the base model, utilizing the pre-trained resources without training from scratch. However, $\operatorname{M_{0}}$ exhibits limited clinical text comprehension. Hence, in Stage 1 we improve the clinical text understanding of $\operatorname{M_{0}}$’s encoder, $\operatorname{ENC_{0}}$, see Figure LABEL:fig:framework. Specifically, we employ $\operatorname{ENC_{0}}$ from $\operatorname{M_{0}}$ and the frozen decoder $\operatorname{DEC_{0}}$ from the original pre-trained model, training $\operatorname{ENC_{0}}$ by reconstructing $N_{i}$. This yields $\operatorname{M}$ with an improved encoder $\operatorname{ENC}$ which better grasps the complex semantics of nursing notes.
@@ 3.2 Training Objective
Since there is no ground truth summary available, in Stage 2, we adopt a self-supervised strategy for model $\operatorname{M}$ to generate high-quality, patient-centered summaries that can respond to patient-related queries effectively (Figure LABEL:fig:framework). We introduce a separate model $\operatorname{R}$, a query responder, trained to handle patient-related queries such as predicting readmission risk using data available in the patient database. For example, given the note $N_{i}$ or its summary $S_{i}$ as input to $\operatorname{R}$, the output will be the response to the corresponding query, like the readmission probability. As the training objective for $\operatorname{M}$ we minimize the discrepancy, calculated as the cross-entropy loss $\mathcal{L}_{CE}$, between the two responses based on either $N_{i}$ or $S_{i}$ respectively. This ensures that when responding to a certain patient-related query, using the summary will produce a response similar to that obtained using the nursing note, without ever showing the correct query answers to model $\operatorname{M}$. To prevent $\operatorname{M}$ from generating summaries that are too verbose or direct “copy-paste” from the original notes, we introduce a length penalty term. The final loss function for a batch of nursing notes is:
$\mathcal{L}_{summ}$ $=\frac{1}{K}\sum_{i=1}^{K}\mathcal{L}_{CE}(\operatorname{R}(N_{i}),\operatorname{R}(S_{i}))$
×$\times(1+\lambda_{1}e^{(\alpha-0.5)}),$ (1)
@@
$\alpha=\frac{\sum_{i=1}^{K}\operatorname{Len}(S_{i})}{\sum_{i=1}^{K}\operatorname{Len}(N_{i})},$ (2)
$K$ is the batch size, and $\operatorname{Len}$ denotes the number of tokens in the document. The hyperparameter $\lambda_{1}\in[0,1]$ regulates the extent of the penalty, and is selected using the validation set, see details in C.2. Nondifferentiability from $\operatorname{M}$ and $\operatorname{R}$ is resolved using the straight-through Gumbel-softmax trick (Bengio et al., 2013; Jang et al., 2017).
@@ 3.3 Augmentation Blocks for the Context of the Patient
A patient typically has sequential nursing notes to document the evolution of their condition. Therefore, the context provided by the prior notes helps understanding the patient’s current status. We regard this as temporal information and include it during summarization to help the model to understand the patient’s condition. For $N_{i}$, the embeddings of its previous notes are represented by the embeddings of their respective first tokens, which are special tokens indicating the start of each note. These embeddings are obtained at the last hidden state in the $\operatorname{ENC}$, denoted as $\{\mathbf{h_{1}},\mathbf{h_{2}},\ldots,\mathbf{h_{i-1}}\}$, where $\mathbf{h_{i}}\in\mathbb{R}^{d}$ and $d$ is the hidden state dimension. We aggregate past note representations via weighted mean pooling, emphasizing the most recent notes. In practice, initial weights $\beta_{j},j=1,\ldots,i-1$ for each past note $N_{i}$ are based on the position in the sequence: $\beta_{1}=1,\beta_{2}=2,\ldots,\beta_{i-1}=i-1$. We use the normalized weights $\beta^{{}^{\prime}}_{j}=\frac{\beta_{j}}{\beta_{1}+\beta_{2}+...+\beta_{i-1}}$ to perform the weighted mean pooling:
@@
where $\mathbf{h}^{TIF}\in\mathbb{R}^{d}$ represents the information fusion of the past notes. As shown in Figure LABEL:fig:augmentation, we prepend a special token [TI] at the beginning of the decoder input, representing temporal information with embedding $\mathbf{h}^{TIF}$. Consequently, the initial input to the decoder consists of [[TI], [BOS]], where [BOS] indicating the start of generation. [TI] is replaced with the padding token [PAD] for nursing notes without past notes.
The model generates tokens auto-regressively, appending produced tokens to the decoder input when generating subsequent tokens. Thus, [TI] consistently prompts the model to focus on the patient’s past context throughout the summarization process.
@@
We aim to obtain summaries focusing on the patient’s condition by incorporating patient-level information into the model via cross-attention, facilitating the interaction of information on different levels. A patient’s metadata PA typically comprises basic information, including age, gender, existing diagnoses, and performed procedures. We convert this metadata into patient information in natural language (one example in A.1), then encode it using $\operatorname{ENC}$ to derive patient embedding $\mathbf{H}^{PA}\in\mathbb{R}^{z\times d}$, where $z$ represents the number of tokens in patient information. The source note is also encoded into embedding $\mathbf{H}^{enc}\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of tokens in the note given as input. On the decoder $\operatorname{DEC}$ side, let us assume the tokens input to the decoder at the current timestep are [[TI], [BOS], $y_{1}$,…,$y_{j}$]. Consequently, the hidden representation passed to the $l$th decoder layer is $\mathbf{H}_{l}^{dec}\in\mathbb{R}^{{(j+2)}\times d}$. $\mathbf{H}_{l}^{dec}$ is updated using Multi-Head Self-Attention ($\operatorname{MHSA}$) and Multi-Head Cross-Attention ($\operatorname{MHCA}$(Vaswani et al., 2017) involving $\mathbf{H}^{enc}$ and $\mathbf{H}^{PA}$. This facilitates the fusion of patient and note-level information:
$\mathbf{H}_{l+1}^{dec}$ $=\operatorname{MHCA}(\mathbf{H}^{enc},\operatorname{MHSA}(\mathbf{H}_{l}^{dec}))$
$+\lambda_{2}\times\operatorname{MHCA}(\mathbf{H}^{PA},\operatorname{MHSA}(\mathbf{H}_{l}^{dec})),$ (4)
@@
$\lambda_{2}\in[0,1]$ is a hyperparameter to control the importance of patient meta information, selected using the validation set, see details in C.3. $\mathbf{H}_{l+1}^{dec}$ is the input to the next decoder layer, or if the $l$th layer is the final layer, it is the input to the language modeling head.